# Multimodal Fusion Balancing Through Game-Theoretic Regularization

### v2

@@ 3 Multimodal Competition Regularizer
$I(X_{1};X_{2};Y)=\underbrace{I(X_{1};Y\mid X_{2})+I(X_{2};Y\mid X_{1})}_{\begin{subarray}{c}\text{Task-Relevant Unique Information}\\ \text{of each modality}\sim\mathcal{L}_{\operatorname{MIPD}}\end{subarray}}+\underbrace{I(X_{1};X_{2})}_{\begin{subarray}{c}\text{Shared Information}\\ \mathcal{L}_{\operatorname{Con}}\end{subarray}}-\underbrace{I(X_{1};X_{2}\mid Y)}_{\begin{subarray}{c}\text{Task-Irrelevant}\\ \text{Shared Information}\sim\mathcal{L}_{\operatorname{CEB}}\end{subarray}}.$ (5)
This decomposition is illustrated by the Venn diagram in Figure 1. The first two terms, the $\operatorname{CMIs}$ $I(X_{1};Y\mid X_{2})$ and $I(X_{2};Y\mid X_{1})$, represent the unique information each modality holds about the target that is not shared with the other modality. Maximizing these terms encourages the model to extract modality-specific task-relevant features. The third term, $I(X_{1};X_{2})$, quantifies shared information between the modalities. Maximizing it encourages the model to align the representations and effectively utilize the common information between [22; 39; 44].
The third term, $I(X_{1};X_{2})$, quantifies how much information the two modalities share. Maximizing this term encourages the model to align the representations and effectively utilize the common information between them [22; 39; 44]. The third term, $I(X_{1};X_{2})$, quantifies the shared information between the modalities. Maximizing this term encourages to explore common information and aligns the representations, similar to how contrastive learning exploits multi-view redundancy to enhance representation quality [22; 39; 44]. The final term, $I(X_{1};X_{2}\mid Y)$, represents task-irrelevant shared information. It penalizes the shared information between modalities but not useful for predicting the target. For each of these terms, we introduce a corresponding loss component. The total regularization loss is defined as follows:
@@
The proposed regularizer consists of three key losses: $\mathcal{L}_{\operatorname{Con}}$, a contrastive loss that captures shared information between modalities, $\mathcal{L}_{\operatorname{CEB}}$, which filters out task-irrelevant information, focusing on relevant features for the downstream task, and $\mathcal{L}_{\operatorname{MIPD}}$, which measures each modality’s affection to the multimodal prediction by assessing output variations under perturbations. The latter has contradicting terms since increasing the importance of one modality decreases the importance of the other, and hence we model it as a game where each modality encoder (with parameters $\theta_{1},\theta_{2}$) competes to increase the importance of its own modality while decreasing the other. In the remainder of this section, we will examine each component individually and further elaborate on framing multimodal competition using game theory.
3.1 Approximating MI Terms
To approximate each $\operatorname{CMI}$ and capture the unique contribution of each modality, the Mutual Information Perturbed Difference ($\operatorname{MIPD}$) serves as a surrogate function, measuring how input perturbations affect the model’s output. By comparing predictions with and without these perturbations, $\operatorname{MIPD}$ estimates how much information each modality provides. If a modality is crucial, altering its input should significantly change the output, revealing its importance.
The perturbed version of modality $X_{1}$ is denoted as $\tilde{X}{1}$. We use permutations as our perturbation method, which will be detailed further in Section 3.3.
Estimating the $\operatorname{CMI}$ directly $I(X_{1};Y\mid X_{2})=H(Y\mid X_{2})-H(Y\mid X_{1},X_{2})$ is typically intractable. Instead we use the $\operatorname{MIPD}$, which is defined as:
$\operatorname{MIPD}(X_{1};Y\mid X_{2})$ $=I(X_{1};Y\mid X_{2})-I(\tilde{X}_{1};Y\mid X_{2})\leq I(X_{1};Y\mid X_{2}).$ (8)
Using the entropy interpretation of the mutual information terms, each $\operatorname{CMI}$ can be expressed as the difference of the $\log$ probabilities with and without the perturbations:
@@ 3.1 Approximating MI Terms
Moreover, we employ the Jensen-Shannon divergence ($\operatorname{JSD}$) [32], which is symmetrically bounded, to avoid training instabilities leading to the following objective:
@@
Next, the $I(X_{1};X_{2})$ $\operatorname{MI}$ term measures how much information the two modalities share with each other. It the common patterns between the two, ensuring that the model recognizes and aligns these shared aspects. We exploit the available label information employing the supervised contrastive loss [24] $\mathcal{L}_{\operatorname{Con}}$. The term is defined as follows:
@@
$I(X_{2};Y|X_{1})+I(X_{1};Y|X_{2})+2\cdot I(X_{2};X_{1})\geq\log N-\mathcal{L}_{\operatorname{Con}}^{Opt}$ (17)
As $N$ increases the bounds becomes tighter, while the bound is not affected by the number of positive samples (same class datapoints). More details are provided in Appendix B.
Finally, $I(X_{1};X_{2}\mid Y)$ captures irrelevant shared information between modalities, and minimizing an upper bound on this ensures the model retains only task-relevant content. For this purpose, we exploit $\operatorname{CEB}$ [7], targeting superfluous information in multimodal representations via a reconstruction loss. A small reconstruction head, $h:Y;\theta_{h}\rightarrow Z=(Z_{1},Z_{2})$, predicts back into the latent space, effectively filtering out irrelevant content:
$\mathcal{L}_{\operatorname{CEB}}(X_{1};X_{2})=\mathbb{E}_{\hskip 1.0pt\begin{subarray}{c}x_{1},x_{2},y\sim p(x_{1},x_{2},y)\end{subarray}}\left\|[f_{1}(x_{1}),f_{2}(x_{2})]-h(p(y\mid x_{1},x_{2}))\right\|^{2}$ (18)
Penalizing irrelevant information has been shown to enhance calibration and robustness [8], but it must be carefully evaluated, as it can introduce constraints that may hinder overall performance.
Figure 3: In this figure, we highlight a specific part of our training process to demonstrate how different competition strategies between modalities are applied. The gradient multiplier adjusts the video encoder’s behavior toward Audio Importance ($\text{Importance}_{2}$). When $k=1$, the video encoder collaborates to boost $\text{Importance}_{2}$. At $k=0$, it remains neutral. At $k=-1$, it competes by reducing $\text{Importance}_{2}$ to increase its own ($\text{Importance}_{1}$). This is based on the principle that Total Importance ($\text{Importance}_{2}+\text{Importance}_{1}$) remains constant.
3.2 The Game of Multimodal Fusion
In this section, we present the game-theoretic approach used to balance the terms of the proposed multimodal training regularizer. The key insight is that the total contribution from all modalities, expressed through modality importance, remains constant. This implies that when one modality’s contribution increases significantly, it may diminish the contributions of other modalities. This allows to frame multimodal competition as a constant-sum game where each modality acts as a player. The $\mathcal{L}_{\operatorname{MIPD}}$ corresponding to each modality represents the payoff function for that modality, where each modality aims to maximize its own contribution to the overall output. The players take concurrent actions, as each modality is updated during every optimization step. We constrain the players’ actions by enforcing a fixed strategy throughout the entire training process. We explore three strategic actions for the modalities:
  • Collaborative Actions: All modalities work together to increase each other’s contributions. The $\mathcal{L}_{\operatorname{MIPD}}$ terms are applied across all parameters, resulting in a joint minimization, $\min\limits_{\theta}\mathcal{L}_{\operatorname{MIPD}}$
  • Independent Actions: Each modality focuses on maximizing its own contribution by optimizing its respective $\mathcal{L}_{\operatorname{MIPD}}$ term, without regard to the contributions of other modalities, leading to $\min\limits_{\theta_{i}}\mathcal{L}_{\operatorname{MIPD}_{X_{i}}}$.
  • Greedy Actions: Each modality seeks to maximize its own contribution by: 1) minimizing its own $\mathcal{L}_{\operatorname{MIPD}}$ term, and 2) maximizing the $\mathcal{L}_{\operatorname{MIPD}}$ terms of other modalities. In the constant-sum game, reducing the influence of other modalities enhances its own contribution, resulting in a min-max strategy, $\min\limits_{\theta_{i}}\max\limits_{\theta_{\neg i}}\mathcal{L}_{\operatorname{MIPD}_{X_{i}}}$  [^1] The notation $\neg i$ refers to the rest of the modalities except $i$..
In Figure 3, we provide an example of the choices the video modality can make, illustrating how it can either assist, ignore, or diminish the importance of the audio modality. By default, we adopt the greedy actions strategy, unless otherwise specified.
3.3 Perturbations
To measure the importance of modality $X_{1}$, we observe the difference in its predictive output when presented with $\{\tilde{X}_{1},X_{2}\}$ compared to the normal input $\{X_{1},X_{2}\}$.
Previous works have explored various perturbation techniques. Gat et al. [10] used additive noise to estimate functional entropy by maximizing output variance. Maximizing this variance can increase sensitivity to noise, undermining the goals of Lipschitz smoothness [1] and adversarial robustness [47]. Ji et al. [21] and Liang et al. [31] proposed task-specific augmentations, that may not be available in most tasks and modalities, while their method relies heavily on the success of these augmentations. Li et al. [28] employed zero-masking to estimate Shapley values and trained directly to maximize them. However, Shapley values can produce arbitrary or unreliable estimates, particularly in high-dimensional settings [35]. Additionally, each of these methods perturbs the input, requiring extra forward passes and proportionally increasing computational and memory demands.
To address the limitations of the suggested perturbations and face their computational need, we propose as perturbation to use within-batch permutations $\sigma_{e}\sim\text{Uniform}(\mathcal{P})$, where $\mathcal{P}$ represents the set of all possible permutations [11]. Permutations differ from previous methods by applying the transformation directly to the latent space, eliminating the need for multiple forward passes through the unimodal encoders, thereby reducing both computational and memory overhead. To analyze the effects of permutations, we examine two cases separately: when the label of the original data matches the label of the permuted pair, and when it does not:
  1. 1.
    Matching Labels: The normal and perturbed latent vectors represent the same class, allowing us to test if changes in task-irrelevant information impact the output, resembling the optimal augmentation described in Liang et al. [31]. Maximizing their output difference can help the model spread the datapoint representations within the class boundaries.
  2. 2.
    Non-Matching Labels: In this case, the normal and perturbed latent vectors belong to different classes. Maximizing their output difference can promote better class discrimination, refining the decision boundaries and reinforcing model’s robustness and calibration on out-of-distribution cases.
A detailed analysis of the computational benefits of applying perturbations to the latent space is provided in Appendix A.6, along with a comparison of perturbations in Appendix C.1.
Algorithm 1 Multimodal Training with MCR
Input: Training dataset $D$ with modalities $X_{1},X_{2},\ldots,X_{n}$, output labels $Y_{t}$, multimodal model $f\in\mathcal{F}$, initialized unimodal encoders $\theta_{i}$, reconstruction model $h$, $\lambda_{\text{uni}},\lambda_{\text{M}}$ Lagrangian coefficients
1:  for each epoch do
2:     for each batch $(X_{1},..,X_{n},Y_{t})$ sampled from $D$ do
3:        Estimate $Y,Z_{1},..,Z_{n}=f(X_{1},..,X_{n})$
4:        Compute $L_{task}(f(X_{1},..,X_{n}),Y_{t})$ and $L_{task}^{uni}=\lambda_{\text{uni}}\sum_{i=1}^{M}L_{task}(f_{i}^{u}(X_{i}),Y_{t})$
5:        Compute the $L_{\text{Con}}$ with Eq. 16 and $L_{\text{CEB}}$ with Eq. 18
6:        Sample $P_{e}$ permutations and compute permuted pairs on the latent space $Z$
7:        Pass each pair through the fusion model $f_{c}$ to get predictions $\tilde{Y}_{i}$ with modality $i$ permuted
8:        Compute $L_{\text{MIPD}}=\lambda_{\text{M}}\sum_{i=1}^{M}L_{\text{MIPD}_{x_{i}}}$ using Eq. 12 and the corresponding strategy
9:        Compute $L_{\operatorname{MCR}}$ from Eq. 6
10:        Update with $L_{total}=L_{task}+L_{uni}+L_{\operatorname{MCR}}$
11:     end for
12:  end for
4 Experiments
We perform extensive experiments on synthetic and real-world multimodal datasets to evaluate the effectiveness of our proposed method, MCR, against various baselines and previous methods:
Unimodal Training: Separate training for each modality.
Ensemble: Combines unimodal predictions without any further training.
Joint Training: Trains all modalities under a single-loss objective, without any balancing.
Multi-Loss [52]: Incorporates additional unimodal task losses alongside the joint objective.
MMCosine [59]: Equalizes modality contributions by standardizing features and weights.
MSLR [60]: Adapts unimodal encoders’ learning rates based on validation performance.
PMR [6]: Similar to Multi-Loss, uses prototype classifiers and adjusts unimodal loss weights.
OGM [40]: Extracts unimodal performance from the last layer, compares improvements, and modulates gradients by adjusting each encoder’s learning rate.
AGM [28]: Uses zero-masking Shapley values as predictor for the unimodal task losses and modulate them by adjusting their coefficients.
MLB [26]: Combines Multi-Loss for enhanced unimodal performance assessment and adjusts each encoder’s learning rate.
Uni-Pre Frozen: Uses pre-trained unimodal encoders with frozen weights during joint training.
Uni-Pre Finetuned: Employs pre-trained unimodal encoders, fine-tuned during joint training.
4.1 Synthetic Dataset
We generate an artificial scenario where the mutual information between the label and the modalities varies, demonstrating the phenomenon of modality competition. Although many factors can contribute to this effect, a comprehensive exploration is beyond the scope of this work. Instead, we focus on the imbalance in modality informativeness to highlight the motivation behind our approach.
Data: We generate task-irrelevant information for each modality by sampling $N_{1},N_{2}\sim\mathcal{N}(0,I)$ and the 5-class label $Y$ from a uniform distribution $Y\sim\text{Uniform}(5)$. Using fixed transformations, similar to Liang et al.[31], each modality is converted into a high-dimensional vector. We relate both modalities to the label through a linear relationship: $X_{1}=N_{1}+Y$ and $X_{2}=N_{2}+Y$. Data points are distributed such that either both modalities contain label information (Shared Information) or only one of the modalities (Unique Information). We vary the percentage of data points with shared and unique information to analyze how models perform under different conditions.
Results: Figure 1 shows the performance on synthetic data, comparing our method ($\operatorname{MCR}$) with several baselines. As the shared information $S$ among the modalities decreases, and the unique information of one modality $U_{1}$ increases while the $U_{2}$ remains constant, we observe for all methods a performance drop. $\operatorname{MCR}$ maintains the highest accuracy across all combinations, demonstrating the slowest decline and highlighting its robustness to such imbalance.
4.2 Real-World Datasets
Datasets: We explore several real-world datasets, primarily with video, optical-flow, audio, and text modalities, that have either been identified as exhibiting significant imbalance among modalities or serve as standard multimodal benchmarks:
CREMA-D [4]: An emotion recognition dataset featuring 91 actors expressing 6 distinct emotions.
AVE [49]: A collection of videos with temporally aligned audio-visual events across 28 categories.
UCF [45]: An action recognition dataset consisting of real-life YouTube videos.
CMU-MOSEI [63]: Large-scale multimodal sentiment analysis dataset with 23k monologue clips.
CMU-MOSI [61]: Multimodal sentiment analysis dataset with over 2k YouTube video clips.
Something-Something (V2) [15]: Contains 220k clips of individuals performing 174 hand actions.
Models: We employ a variety of models and backbone encoders to examine the behavior of both smaller-scale models trained from scratch and larger, more complex models pretrained with self-supervised learning (SSL). This combination allows us to demonstrate that our method is effective in both limited data scenarios without pretraining and in cases with ample data where the goal is fine-tuning. We utilize ResNet-18 and small-scale Transformers as alongside with state-of-the-art models like Conformer [14] Swin-TF [33]. Detailed model configurations for each dataset are provided in Appendix A.2. These choices aim to bridge the gap between theoretical work and practical application.
Table 1: This table compares the proposed MCR method with several baselines across the additional datasets: CMU-MOSEI, CMU-MOSI, and Something-Something. Results are shown for different modality combinations (e.g., V-A-T: Video-Audio-Text), using Transformer and Swin-TF backbones. MCR consistently outperforms other methods across all configurations.
Transformer Swin-TF
MOSEI MOSI Something-Something
Method V-T V-A-T V-T V-A-T V-OF
Unimodals V: 63.6 A: 64.3 T: 79.9 V: 53.5 A: 54.9 T: 70.7 V:61.6 OF:50.6
Ensemble 77.4 76.6 69.8 43.3 64.8
Joint Training 80.5 81.2 74.7 73.9 56.3
Multi-Loss 80.8 79.9 75.3 72.0 59.6
OGM [40] 80.6 - 75.2 - 58.3
AGM [28] 79.5 80.6 75.3 75.2 56.9
MLB [26] 80.8 81.1 75.0 74.5 61.3
Uni-Pre Frozen 80.5 79.8 70.6 70.6 64.2
Uni-Pre Finetuned 80.7 80.1 73.8 71.2 61.9
@@ 4.2 Real-World Datasets
Results: In Figure 4 and in Table 1 we present the accuracy comparison among the different baseline methods on our datasets. In the figure we provide accuracy as a comparison with that of Ensemble to highlight the difference between multimodal and unimodal training capabilities. We use both backbone encoders where we identified good SSL pretrained fit models: Conformer [14](in orange) and ResNet (in blue). Positive values indicate an improvement over the ensemble baseline, while negative values signify a decline in performance. Our main observations are:
@@
    Most previous methods and baselines fail to outperform the Ensemble, often showing negative accuracy differences, particularly on the CREMA-D and AVE datasets. These methods do not adequately address the modality competition problem, leading to a worse generalization error struggling to leverage the given multimodal data, regardless of the chosen backbone model.

### v3

@@ 3 Multimodal Competition Regularizer
$I(X_{1};X_{2};Y)=$ $\underbrace{I(X_{1};Y\mid X_{2})+I(X_{2};Y\mid X_{1})}_{\begin{subarray}{c}\text{Task-Relevant Unique Information}\\ \text{of each modality}\sim\mathcal{L}_{\operatorname{MIPD}}\end{subarray}}+\hskip-3.0pt\underbrace{I(X_{1};X_{2})}_{\begin{subarray}{c}\text{Shared Information}\\ \mathcal{L}_{\operatorname{Con}}\end{subarray}}\hskip-3.0pt-\hskip-8.0pt\underbrace{I(X_{1};X_{2}\mid Y)}_{\begin{subarray}{c}\text{Task-Irrelevant}\\ \text{Shared Information}\sim\mathcal{L}_{\operatorname{CEB}}\end{subarray}}\hskip-10.0pt.$ (4)
This decomposition is illustrated by the Venn diagram in Figure 1. The $\operatorname{CMIs}$ $I(X_{1};Y\mid X_{2})$ and $I(X_{2};Y\mid X_{1})$ capture modality-specific information for predicting the target. Maximizing them with the Mutual Information Perturbed Difference ($\operatorname{MIPD}$) loss, $\mathcal{L}_{\operatorname{MIPD}}$, which assesses each modality’s contribution via output variations under input perturbations (elaborated in Sec. 3.3) and encourages the extraction of modality-specific, task-relevant features. The third term, $I(X_{1};X_{2})$, quantifies shared information between modalities. Maximizing it with a contrastive loss, $\mathcal{L}_{\operatorname{Con}}$, aligns representations and leverages their shared information effectively [42, 43, 7]. The final term, $I(X_{1};X_{2}\mid Y)$, represents task-irrelevant shared information. Penalizing it with the conditional entropy bottleneck ($\operatorname{CEB}$) [44] and the corresponding loss $\mathcal{L}_{\operatorname{CEB}}$ to filter out irrelevant information, focusing the model on features relevant to the downstream task. Each term has a corresponding loss, as illustrated in Figure 2, forming the regularizer with three key losses:
$\mathcal{L}_{\operatorname{MCR}}=\mathcal{L}_{\operatorname{MIPD}}+\mathcal{L}_{\operatorname{Con}}+\mathcal{L}_{\operatorname{CEB}}$
Figure 3: This figure illustrates a key aspect of our training process, showing how competition strategies between modalities are applied. The gradient multiplier adjusts the video encoder’s response to Audio Importance ($\text{Importance}_{2}$). When $k=1$, the video encoder enhances $\text{Importance}_{2}$; at $k=0$, it remains neutral, and at k=1k=-1, it competes by reducing $\text{Importance}_{2}$ to prioritize its own ($\text{Importance}_{1}$). This reflects the principle that increasing the importance of one modality can reduce the importance of the other.
3.1 Approximating MI Terms
$\mathbf{I(X_{1};Y\mid X_{2}):}$ To approximate each $\operatorname{CMI}$ and capture the unique contribution of each modality, the $\operatorname{MIPD}$ serves as a surrogate function, measuring how input perturbations affect the model’s output. By comparing predictions with and without these perturbations, $\operatorname{MIPD}$ estimates how much information each modality provides. If a modality is crucial, altering its input should significantly change the output, revealing its importance.
Estimating the $\operatorname{CMI}$ directly through $I(X_{1};Y\mid X_{2})=H(Y\mid X_{2})-H(Y\mid X_{1},X_{2})$ is typically intractable. Instead we use the $\operatorname{MIPD}$ as a lower bound, defined as:
$\operatorname{MIPD}(X_{1};Y\mid X_{2})$ $=I(X_{1};Y\mid X_{2})-I(\tilde{X}_{1};Y\mid X_{2})\leq I(X_{1};Y\mid X_{2}),$ (6)
where the perturbed version of modality $X_{1}$ is denoted as $\tilde{X}_{1}$. Interpreting MI via entropy, each $\operatorname{CMI}$ can be expressed as the difference of the $\log$ probabilities with and without the perturbations:
$=\hskip-28.0pt\underset{\raisebox{-7.0pt}{$\hskip 10.0pt\scriptstyle\begin{subarray}{c}y\sim p(y)\\ \hskip 7.0ptx_{1},x_{2}\sim p(x_{1},x_{2},y)\end{subarray}$}}{\mathbb{E}}\hskip-23.0pt\left[-\hskip-14.0pt\underset{\raisebox{-2.0pt}{$\hskip 10.0pt\scriptstyle\begin{subarray}{c}\tilde{x}_{1}\sim p(x_{1})\end{subarray}$}}{\mathbb{E}}\hskip-12.0pt\left[\log p(y\mid x_{2},\tilde{x}_{1})\right]+\log p(y\mid x_{2},x_{1})\right].$ (7)
Instead of this log-likelihood ratio, we use the symmetrically bounded Jensen-Shannon divergence ($\operatorname{JSD}$) [45] to prevent training instabilities, leading to the following:
$\begin{split}\mathcal{L}_{\operatorname{MIPD_{1}}}&=-\operatorname{MIPD}(X_{1};Y\mid X_{2})=-\hskip-30.0pt\underset{\raisebox{-10.0pt}{$\hskip 10.0pt\scriptstyle\begin{subarray}{c}y\sim p(y)\\ \hskip 7.0ptx_{1},x_{2}\sim p(x_{1},x_{2},y)\\ \tilde{x}_{1}\sim p(x_{1})\end{subarray}$}}{\mathbb{E}}\hskip-25.0pt\left[\operatorname{JSD}(p(y\mid x_{2},x_{1}),p(y\mid x_{2},\tilde{x}_{1}))\right].\end{split}$ (8)
Similarly, $\mathcal{L}_{\operatorname{MIPD}_{2}}$ can be computed symmetrically.
$\mathbf{I(X_{1};X_{2}):}$ The next $\operatorname{MI}$ term measures how much information the two modalities share, capturing the common patterns between the modalities and aligning the representations of these shared aspects. We exploit the available label information employing the supervised contrastive loss $\mathcal{L}_{\operatorname{Con}}$ [46]:
[…]
$I(X_{2};Y|X_{1})+I(X_{1};Y|X_{2})+2I(X_{2};X_{1})\geq\log N-\mathcal{L}_{\operatorname{Con}}^{Opt}.$ (10)
As $N$ increases, the bound becomes tighter, while the bound is not affected by the number of positive samples (same class datapoints). More details are provided in Appendix B.
$\mathbf{I(X_{1};X_{2}\mid Y):}$ The final term captures irrelevant shared information between modalities, and minimizing an upper bound on this ensures the model retains only task-relevant content. For this purpose, we exploit the idea of Conditional Entropy Bottleneck ($\operatorname{CEB}$) $L_{\operatorname{CEB}}$ [47], targeting superfluous information in multimodal representations via a reconstruction loss. A small reconstruction head, $h:Y;\theta_{h}\rightarrow Z=(Z_{1},Z_{2})$, predicts back the latent space, effectively filtering out irrelevant content:
$\mathcal{L}_{\operatorname{CEB}}=\hskip-35.0pt\underset{\raisebox{-5.0pt}{$\hskip 10.0pt\scriptstyle\begin{subarray}{c}x_{1},x_{2},y\sim p(x_{1},x_{2},y)\end{subarray}$}}{\mathbb{E}}\hskip-35.0pt\left\|[f_{1}(x_{1}),f_{2}(x_{2})]-h(y;\theta_{h})\right\|^{2}$ (11)
The exact derivation of this loss term can be found in Appendix A.6. Penalizing irrelevant information has been shown to enhance calibration and robustness [44], but it must be carefully evaluated, as it can introduce constraints that may hinder overall performance.
3.2 The Game of Multimodal Fusion
We adopt a game-theoretic approach to balance the terms of the proposed $\mathcal{L}_{\operatorname{MIPD}}$. The key idea is that increasing one modality’s importance (e.g., via $\operatorname{MIPD}_{1}$) can inherently reduce the other’s (e.g., $\operatorname{MIPD}_{2}$). Thus, an underutilized encoder $i$ (with parameters $\theta_{i}$) can boost its relevance both by minimizing $\mathcal{L}_{\operatorname{MIPD}_{i}}$ and by maximizing $\mathcal{L}_{\operatorname{MIPD}_{\neg i}}$. This twofold strategy helps prevent suppression of weaker modalities. We frame $\mathcal{L}_{\operatorname{MIPD}}$ as a game where each encoder (player) selects a strategy, minimize, maximize, or ignore. Figure 3 illustrates how the video modality, via a hyperparameter $k$, can choose to assist, ignore, or diminish the audio modality. Each encoder applies this logic selectively as formalized below:
$\nabla_{\theta_{1}}\mathcal{L}_{\operatorname{MIPD}}$ $=\lambda_{M}\left(\nabla_{\theta_{1}}\mathcal{L}_{\operatorname{MIPD}_{1}}+k\,\nabla_{\theta_{1}}\mathcal{L}_{\operatorname{MIPD}_{2}}\right),$ (12)
$\nabla_{\theta_{2}}\mathcal{L}_{\operatorname{MIPD}}$ $=\lambda_{M}\left(\nabla_{\theta_{2}}\mathcal{L}_{\operatorname{MIPD}_{2}}+k\,\nabla_{\theta_{2}}\mathcal{L}_{\operatorname{MIPD}_{1}}\right).$ (13)
where $\lambda_{M}$ is a Lagrange multiplier, and $k\in\{-1,0,1\}$ sets the modality’s strategy:
  • Collaborative ($k=1$): All modalities work together to increase each other’s contributions. The $\mathcal{L}_{\operatorname{MIPD}}$ terms are applied across all parameters, resulting in $\min\limits_{\theta}\mathcal{L}_{\operatorname{MIPD}}$.
  • Independent ($k=0$): Each modality focuses on maximizing its own contribution by optimizing solely its respective $\mathcal{L}_{\operatorname{MIPD}}$ term, leading to $\min\limits_{\theta_{i}}\mathcal{L}_{\operatorname{MIPD}_{i}}$.
  • Greedy (k=1k=-1): Each modality seeks to maximize its own contribution by: 1) minimizing its own $\mathcal{L}_{\operatorname{MIPD}}$ term, and 2) maximizing the $\mathcal{L}_{\operatorname{MIPD}}$ terms of other modalities, resulting in a min-max game, $\min\limits_{\theta_{i}}\max\limits_{\theta_{\neg i}}\mathcal{L}_{\operatorname{MIPD}_{i}}$ [^3] The notation $\neg i$ refers to the rest of the modalities except $i$..
Following the results in Appendix A.8, we adopt the greedy strategy as default, as it showed the most consistent performance in our setting.
3.3 Perturbations
To assess the importance of modality $X_{1}$, we define $\mathcal{L}_{\operatorname{MIPD_{1}}}$, which captures changes in the model’s output when $X_{1}$ is perturbed (i.e., $\{\tilde{X}_{1},X_{2}\}$ vs. $\{X_{1},X_{2}\}$). Instead of traditional input-space perturbations, which can be computationally expensive and task-dependent, we apply a within-batch permutation $\sigma_{e}\sim\text{Uniform}(\mathcal{P})$ in the latent space, yielding $\tilde{X}_{1}=\sigma_{e}(X_{1})$. This approach avoids extra forward passes and reduces computational and memory overhead. Further analysis of this technique and comparisons with prior methods are provided in Appendix A.10 and A.14.
The complete algorithm is presented in Algorithm 1, with an extension of $\mathcal{L}_{\operatorname{MCR}}$ to $M$ modalities described in Appendix A.7. In Appendix A.9, we analyze various combinations of loss components, revealing that penalizing task-irrelevant information benefits models with extensive SSL pretraining but proves detrimental for those without it.
Algorithm 1 Multimodal Training with MCR
Input: Training dataset $D$ with modalities $X_{1},X_{2},\ldots,X_{M}$, labels $Y_{t}$, multimodal model $f\in\mathcal{F}$, initialized unimodal encoders $\theta_{i}$, reconstruction model $h$, $\lambda_{\text{uni}},\lambda_{\text{M}}$ Lagrangian coefficients:
1:for each batch $(X_{1},..,X_{M},Y_{t})$ of each epoch do
2:  Compute $\mathcal{L}_{task}(f(X_{1},..,X_{M}),Y_{t})$ and $\mathcal{L}_{task}^{uni}=\lambda_{\text{uni}}\sum_{m=1}^{M}\mathcal{L}_{task}(f_{m}^{u}(X_{m}),Y_{t})$
3:  Extract $(Z_{1},..,Z_{M})$ from $f(X_{1},..,X_{M})$
4:  Assess the $\mathcal{L}_{\text{Con}}$ with Eq. 9 and $\mathcal{L}_{\text{CEB}}$ with Eq. 11
5:  Sample $\sigma_{e}$ permutations and compute permuted pairs on the latent space $Z$
6:  Pass each pair through the fusion model $f_{c}$ to get predictions $\tilde{Y}_{m}$ with modality $m$ permuted
7:  Compute $\mathcal{L}_{\text{MIPD}}$ using Eq. 13, 8 and selecting $k$ by strategy (default: Greedy)
8:  Determine $\mathcal{L}_{\operatorname{MCR}}$ from Eq. 5
9:  Update the model parameters based on $\mathcal{L}_{total}=\mathcal{L}_{task}+\mathcal{L}_{uni}+\mathcal{L}_{\operatorname{MCR}}$
10:end for
4 Experiments
4.1 Synthetic Dataset
We create a scenario where mutual information varies, showcasing modality competition. While various factors can contribute to such a phenomenon, we focus on modality informativeness imbalance to motivate our approach.
Data: We generate task-irrelevant information for each modality by sampling $N_{1},N_{2}\sim\mathcal{N}(0,\mathbf{I})$ and the 5-class label $Y_{t}$ from a uniform distribution $Y_{t}\sim\text{Uniform}(5)$. Each modality is converted into a high-dimensional vector using fixed transformations, similar to Liang et al.[41]. We relate both modalities to the label through a linear relationship: $X_{1}=N_{1}+Y_{t}$ and $X_{2}=N_{2}+Y_{t}$. Data points are distributed in such a way that either both modalities contain label information (Shared Information) or only one of the modalities (Unique Information). In cases where only one modality contains label information, the other modality is defined as $X_{1}=N_{1}$ and $X_{2}=N_{2}$ respectively. We vary the percentage of data points with shared and unique information to analyze model performance under different conditions.
Results: Figure 1 shows the performance on synthetic data, comparing our method ($\operatorname{MCR}$) with several baselines. As the shared information $S$ among the modalities decreases, and the unique information of one modality $U_{1}$ increases while the $U_{2}$ remains constant, we observe a performance drop for all methods. $\operatorname{MCR}$ maintains the highest accuracy across all combinations, demonstrating the slowest decline and highlighting its robustness to such imbalance.
4.2 Real-World Datasets
Datasets: We explore several real-world datasets, primarily with video, optical flow, audio, and text modalities, that either exhibit significant imbalance among modalities or serve as standard multimodal benchmarks. Detailed descriptions are provided in Appendix A.3, with brief summaries below:
1. CREMA-D [32]: An emotion recognition dataset with 91 actors expressing 6 distinct emotions.
2. AVE [30]: A collection of videos with temporally aligned audio-visual events across 28 categories.
3. UCF [31]: An action recognition dataset of real-life YouTube videos.
4. CMU-MOSEI [34]: Multimodal sentiment analysis dataset with 23k monologue clips.
5. CMU-MOSI [33]: Multimodal sentiment analysis dataset with over 2k YouTube video clips.
6. Something-Something (V2) [35]: 220k clips of individuals performing 174 hand actions.
Models: We employ a variety of models and backbone encoders to examine the behavior of both smaller-scale models trained from scratch and larger, more complex models pretrained with self-supervised learning (SSL). This combination demonstrates that our method is effective in both limited data scenarios without pretraining and in cases with ample data where the goal is fine-tuning. We utilize ResNet-18 and small-scale Transformers (from thousand to 20M parameters) alongside state-of-the-art models such as Swin-TF [48] and Conformer [49], incorporating backbone encoders like Wav2Vec2, HuBERT, and ViViT, resulting in model sizes approaching 200M parameters. Detailed model configurations for each dataset are provided in Appendix A.4 and experimental details in Appendix A.5. These choices aim to bridge the gap between theoretical work and practical application.
Table 1: Performance comparison of $\operatorname{MCR}$ against prior multimodal training methods across six datasets. $\operatorname{MCR}$ consistently achieves top results across various modality combinations (V-A, V-T, V-A-T, V-OF) and model architectures. MOSI and MOSEI use three modalities and are trained as regression tasks (converted to binary accuracy); the rest are classification tasks. All baselines were rerun under our evaluation protocol to ensure fair comparison and address data leakage issues identified in previous evaluation setups.
CREMA-D AVE UCF MOSI MOSEI Sth-Sth
ResNet Conformer ResNet Conformer ResNet Transformer Swin-TF
Method V-A V-A V-A V-A V-A V-T V-A-T V-T V-A-T V-OF
Unimodals
@@ 4.2 Real-World Datasets
[…]
$\times$ method not applicable to regression tasks;
$\otimes$ method not applicable to trimodal inputs;
- result not reported.
Results: Table 1 reports accuracy comparisons across baseline methods on our evaluation datasets. We highlight two key observations:
@@
    Most prior methods, including recent multimodal approaches, fail to outperform simpler alternatives such as Ensembles or unimodal encoders (either frozen or finetuned). While some of these methods partially address modality competition, they often fall short of effectively leveraging multimodal data.